A TMS (transcranial magnetic stimulation) device creates very rapidly changing magnetic fields. The field near a typical pulsed-field machine rises from0 T to 2.5 T in 200 μs. Suppose a technician holds his hand near the device so that the axis of his 2.3-cm-diameter wedding band is parallel to the field.

What emf is induced in the ring as the field changes?

If the band is made of a gold alloy with resistivity 6.2×10−8Ω⋅m and has a cross-section area4.5 mm2 , what is the induced current?
Answer #1

Induced emf = - A dB / dt

Induced emf = E = - A ( B2 - B1)/t

Here

A = 0.25 pi d^2 = 0.25 * pi * 0.023^2 =

B2 = 2.5 T

B1 = 0 T

t = 200 us = 200 * 10^-6 s

E = -  0.25 * pi * 0.023^2 * (2.5 - 0)/(200*10^-6) = - 5.193 V

(a) Induced emf = 5.913 V

(b) Induced current = E / R = 5.913 / ( 6.2 * 10^-8 * pi * 0.023 / (4.5*10^-4)) = 593951.515 A